OpenStudy (anonymous):

FAN AND MEDAL!!!!! Part A: Divide (10x4y3 + 5x3y2 - 15x2y - 25x2y4) by -5x2y. Show your work, and justify each step. (6 points) Part B: How would your answer in Part A be affected if the x2 variable in the denominator was just an x? (2 points) Part C: What is the degree and classification of the polynomial you got in Part A? (2 points)

OpenStudy (anonymous):

do you know how to do it then?

OpenStudy (anonymous):

ok

OpenStudy (agreene):

\[\frac{10x^4y^3 + 5x^3y^2 - 15x^2y - 25x^2y^4}{-5x^2y}\] is the same as \[-\frac{10x^4y^3}{5x^2y}-\frac{5x^3y^2}{5x^2y}+\frac{15x^2y}{5x^2y}+\frac{25x^2y^4}{5x^2y}\] simplify each of these fractions and you should be on your way.

OpenStudy (anonymous):

Idk how to do that, thats why im here. Im sorry but what you put there was the only obvious part...:/

OpenStudy (anonymous):

you dont need to use the fancy stuff, just using / and ^ works for me and ita alot faster

OpenStudy (agreene):

well, lets take the 1st one: \[-\frac{10x^4y^3}{5x^2y}\] we look at what variables we have on top and bottom, see if we can cancel anything, including any constants: \[-\frac{\color{red}{15}\color{blue}{x^4}y^3}{\color{red}{5}\color{blue}{x^2}y}\]

OpenStudy (agreene):

we can cancel the entire bottom and get: \[-5x^2y^2\] try and do the other ones.

OpenStudy (anonymous):

how did the 10 turn into 15?

OpenStudy (agreene):

cause i wrote it wrong lol

OpenStudy (agreene):

10/5 = 2 so it should be -2x^2y^2

OpenStudy (anonymous):

so the second one would be -x^2y^2?

OpenStudy (agreene):

no, it would be -xy

OpenStudy (agreene):

5/5 = 1 constant 3-2 = 1 x power 2-1 = 1 y power

OpenStudy (anonymous):

so then so far what we would have would be 2x^2y^2 - xy?

OpenStudy (agreene):

right

OpenStudy (agreene):

well, u dropped a - off the front actually

OpenStudy (agreene):

\(-2x^2y^2-xy\) is where we are now

OpenStudy (anonymous):

so then if you add the third part you have -2x^2y^2 - xy + 3x^2y?

OpenStudy (agreene):

no, the 3rd one \[\frac{15x^2y}{5x^2y}=3\]

OpenStudy (anonymous):

then if you add the last one you have -2x^2y^2 - xy + 3 + 5y^4?

OpenStudy (agreene):

5y^3 but yes

OpenStudy (anonymous):

then is that the end or do i need to do something else to it?

OpenStudy (agreene):

well, the normal way to write polynomials is to decent in exponential order and to not lead with a negative, so you would rearrange it to \[5y^3-2x^2y^2-xy+3\]

OpenStudy (anonymous):

oh, that makes sence. how do i know what the degree and classification are?

OpenStudy (agreene):

degree is the number of exponents, in this case: 3+2+2+1+1 = 9

OpenStudy (anonymous):

so then 9th degree? and is it binomal?

OpenStudy (agreene):

the classification comes from the highest exponent, in this case 3, which is known as a cubic

OpenStudy (anonymous):

so trinominal?

OpenStudy (agreene):

yes, but i actually told you wrong on the degree

OpenStudy (agreene):

you look at each term and the one with the highest is the degree of the whole thing

OpenStudy (agreene):

so, its 3 4 2 so its a 4th degree

OpenStudy (anonymous):

thank you so much!

OpenStudy (agreene):

no problem!